# Gender Disparities in Contributions, Leadership, and Collaboration: An Exploratory Study on Software Systems Research

### v1

@@ I Introduction
Several studies have examined female participation in the software industry [14, 15, 16] and software research [17, 18, 19, 20, 21]. For instance, Cavero et al. [21] reported an annual 3.5% growth in women’s participation from 1960 to 2010. However, they focused mainly on the human-computer interaction (HCI) field. Mathew et al. [22] highlighted women’s underrepresentation in software engineering articles. Studies also showed that many women leave scientific careers due to the lack of recognition for their contributions [23, 24]. Therefore, despite several studies on software research focusing primarily on participation statistics, an in-depth analysis of women’s specific contributions, leadership, and collaboration dynamics is warranted.
@@ II-A Dataset Construction
II-B Analyzing the Prevalence of Female Authorship and Their Leadership
To analyze the prevalence of female authors, we first recorded the names of all authors for each article. We determined their genders using the biographies provided at the end of the articles. We then analyzed female authors’ prevalence and their leadership and supervisory roles. Specifically:
@@ II-B Analyzing the Prevalence of Female Authorship and Their Leadership
II-C Analyzing the Contributions of Female Authors
@@
II-D Identifying Female Authors’ Contributions in the Key Areas of Software Systems Research
@@
II-E Exploring the Collaboration Dynamics of Female Authors
@@
III Study Findings
III-A Participation and Leadership of Female Authors (RQ1)
Overall, approximately half of the articles include one or more female authors, while 47.71% have no female authors at all. Furthermore, only 15 out of 763 articles are authored exclusively by females (Fig LABEL:fig:authors_ratio). Although a slight majority of the articles involve female participation, the significant proportion of studies without any female authors highlights a persistent gender imbalance in the field. To further investigate gender representation, we analyzed 384 articles that included both male and female authors. As shown in Fig. LABEL:fig:total_male_female, there is a noticeable underrepresentation of females in software systems research. Among the 1,771 authors across these articles, only 580 (32.7%) are female, while 1,191 (67.3%) are male. This disparity indicates that women may also be underrepresented in influential roles like leadership and supervision.
We then analyzed the prevalence of women in leadership and supervisory roles. As shown in Fig.LABEL:fig:supervised_studies, female-led studies represent an impressive percentage (41.92%), but they remain fewer than male-led studies, which account for 58.07%. Similarly, women are underrepresented in supervisory roles, with their participation nearly 8% lower than that of men. This underrepresentation may be influenced by perceptions, such as the belief that female leaders are less effective than their male counterparts [32], potentially limiting mentorship opportunities and career growth for female researchers.
@@ III-A Participation and Leadership of Female Authors (RQ1)
III-B Commitment and Contributions of Female Authors (RQ2)
The previous section highlighted the overall prevalence of women’s participation in SSR. This section analyzes their involvement across various contributory roles. Fig. 3 illustrates notable patterns in female authors’ involvement in contribution roles. We excluded the funding acquisition and resources roles from the figure as participation from both male and female authors in these roles is minimal for quartile analysis. Specifically, 343 out of 384 studies had no female authors contributing to funding acquisition, and 312 out of 384 studies had no female authors involved in resource-related tasks. Similarly, 278 studies lacked male contributors in funding acquisition, and 297 studies had no male authors in the resources role. This suggests that both male and female authors contributed very little to funding and resource management tasks, with these roles mostly handled by very few individuals considered outliers (Section II-C).
@@ III-B Commitment and Contributions of Female Authors (RQ2)
III-C Leadership and Contribution of Female Authors’ Across Key Areas (RQ3)
@@
III-D Collaboration Dynamics of Female Authors (RQ4)
@@
Next, we investigated the collaboration dynamics when a female author holds the lead or supervisory position and conducted a similar analysis for male authors as well. For investigating the collaboration dynamic in the supervisory positions, we focused on studies where either a female or a male was in a supervisory position (Section, II-B).
@@
IV Key findings
Fading Women, Rising Disparity. Our studies found that over 47% (364 out of 763) of studies do not have any female authors at all (Section III-A). Even in those studies, female participants are significantly low. While female-led studies have a higher ratio (41.92%) than female-supervised studies (25.34%), Fig. 5 shows that female-led studies have been declining over the past five years. After reaching its highest point at 53.33% in 2021, the percentage of female-led studies has steadily declined through 2024. Conversely, female-supervised studies (Fig. 6) have gradually increased, but their ratio is still lower than male-supervised studies. Even when male-supervised studies were at their lowest in 2023, female-supervised studies did not reach a similarly high level. These findings align with research by Kohl et al. [37], which highlights the challenges women face in software engineering management roles. Women in leadership roles often feel they must work harder than their male peers to gain the same level of recognition. There is a perception that when men perform the same work, it is praised, whereas when women do it, it is viewed with question, highlighting the additional pressures women face in these roles.
2020202120222023202430%35%40%45%50%female-ledmale-led
@@ IV Key findings
Women in Software Systems Research: Still Undervalued. Our study found that out of 14 contributory roles, women are involved equally or similarly in most of the roles (Section III-B). Moreover, their contribution is significantly higher in outlining the study objective and preparing the manuscript. However, as shown in Fig. 7, the number of studies where women had no contribution to a specific role is consistently lower than that of men. The lower rate of contribution by women in SSR teams may stem from their undervaluation within these groups. Research indicates that women often face challenges such as gender bias, lack of recognition, and limited access to leadership opportunities, which can hinder their active participation and contribution. For instance, a study by Guzmán [38] highlights that women in the software industry frequently encounter socio-cultural challenges that discourage their involvement, leading to feelings of isolation and undervaluation.
@@ VI Threats to Validity
Threats to external validity relate to the generalizability of our findings. Our study focuses on software systems research using the Journal of Systems and Software (JSS) as a key source. While other journals cover related topics, they often lack a dedicated section highlighting author contributions. By analyzing all JSS articles since the introduction of CRediT, we ensured thorough coverage for this journal. Our findings can be applied to other journals using the CRediT taxonomy but may not generalize to fields with different contribution recording methods.

### v2

@@ I Introduction
Several studies have examined female participation in the software industry [14, 15, 16] and software research [17, 18, 19, 20, 21]. For instance, Cavero et al. [21] reported an annual 3.5% growth in women’s participation from 1960 to 2010. However, they focused mainly on the human-computer interaction (HCI) field. Mathew et al. [22] highlighted women’s underrepresentation in software engineering articles. Studies also show that many women leave scientific careers due to the lack of recognition for their contributions [23, 24]. Therefore, despite several studies on software research focusing primarily on participation statistics, an in-depth analysis of women’s specific contributions, leadership, and collaboration dynamics is warranted.
@@ III-A Dataset Construction
III-B Analyzing the Prevalence of Female Authorship and Their Leadership
To analyze the prevalence of female authors, we first recorded the names of all authors for each article. We identified their genders using the biographies provided at the end of the articles. We then analyzed female authors’ prevalence and their leadership and supervisory roles. Specifically:
@@ III-B Analyzing the Prevalence of Female Authorship and Their Leadership
III-C Analyzing the Contributions of Female Authors
@@
III-D Identifying Female Authors’ Contributions in the Key Areas of Software Systems Research
@@
III-E Exploring the Collaboration Dynamics of Female Authors
@@
IV Study Findings
IV-A Participation and Leadership of Female Authors (RQ1)
Overall, approximately half of the articles include one or more female authors, while 47.71% have no female authors at all. Furthermore, only 15 out of 763 articles are authored exclusively by females (Fig 2a). Although a slight majority of the articles involve female participation, the significant proportion of studies without any female authors highlights a persistent gender imbalance in the field. To further investigate gender representation, we analyzed 384 articles that included both male and female authors. As shown in Fig. 2b, there is a noticeable underrepresentation of females in software systems research. Among the 1,771 authors across these articles, only 580 (32.7%) are female, while 1,191 (67.3%) are male. This disparity indicates that women may also be underrepresented in influential roles like leadership and supervision.
We then analyzed the prevalence of women in leadership and supervisory roles. As shown in Fig. 2c, female-led studies represent an impressive percentage (41.92%), but they remain fewer than male-led studies, which account for 58.07%. Similarly, women are underrepresented in supervisory roles, with their participation nearly 8% lower than that of men. This underrepresentation may be influenced by perceptions, such as the belief that female leaders are less effective than their male counterparts [40], potentially limiting mentorship opportunities and career growth for female researchers.
@@ IV-A Participation and Leadership of Female Authors (RQ1)
IV-B Commitment and Contributions of Female Authors (RQ2)
The previous section highlighted the overall prevalence of women’s participation in SSR. This section analyzes their involvement across various contributory roles. Fig. 3 illustrates notable patterns in female authors’ involvement in contribution roles. We excluded the funding acquisition and resources roles from the figure as participation from both male and female authors in these roles is minimal for quartile analysis. Specifically, 343 out of 384 studies had no female authors contributing to funding acquisition, and 312 out of 384 studies had no female authors involved in resource-related tasks. Similarly, 278 studies lacked male contributors in funding acquisition, and 297 studies had no male authors in the resources role. This suggests that both male and female authors contributed very little to funding and resource management tasks, with these roles mostly handled by very few individuals considered outliers (Section III-C).
@@ IV-B Commitment and Contributions of Female Authors (RQ2)
IV-C Leadership and Contribution of Female Authors’ Across Key Areas (RQ3)
@@
IV-D Collaboration Dynamics of Female Authors (RQ4)
@@
Next, we investigated the collaboration dynamics when a female author holds the lead or supervisory position and conducted a similar analysis for male authors as well. For investigating the collaboration dynamic in the supervisory positions, we focused on studies where either a female or a male was in a supervisory position (Section, III-B).
@@
V Key findings
Fading Women, Rising Disparity. Our studies found that over 47% (364 out of 763) of studies do not have any female authors at all (Section IV-A). Even in those studies, female participants are significantly low. While female-led studies have a higher ratio (41.92%) than female-supervised studies (25.34%), Fig. 5 shows that female-led studies have been declining over the past five years. After reaching its highest point at 53.33% in 2021, the percentage of female-led studies has steadily declined through 2024. Conversely, female-supervised studies (Fig. 6) have gradually increased, but their ratio is still lower than male-supervised studies. Even when male-supervised studies were at their lowest in 2023, female-supervised studies did not reach a similarly high level. These findings align with research by Kohl et al. [45], which highlights the challenges women face in software engineering management roles. Women in leadership roles often feel they must work harder than their male peers to gain the same level of recognition. There is a perception that when men perform the same work, it is praised, whereas when women do it, it is viewed with question, highlighting the additional pressures women face in these roles.
2020202120222023202430%35%40%45%50%femalemale
@@ V Key findings
Women in Software Systems Research: Still Undervalued. Our study found that out of 14 contributory roles, women are involved equally or similarly in most of the roles (Section IV-B). Moreover, their contribution is significantly higher in outlining the study objective and preparing the manuscript. However, as shown in Fig. 7, the number of studies where women had no contribution to a specific role is consistently lower than that of men. The lower rate of contribution by women in SSR teams may stem from their undervaluation within these groups. Research indicates that women often face challenges such as gender bias, lack of recognition, and limited access to leadership opportunities, which can hinder their active participation and contribution. For instance, a study by Guzmán [46] highlights that women in the software industry frequently encounter socio-cultural challenges that discourage their involvement, leading to feelings of isolation and undervaluation.
@@ VI Threats to Validity
Threats to external validity relate to the generalizability of our findings. Our study focuses on software systems research using the JSS as a key source. While other journals cover related topics, they often lack a dedicated section highlighting author contributions. By analyzing all JSS articles since the introduction of CRediT, we ensured thorough coverage for this journal. Our findings can be applied to other journals using the CRediT taxonomy but may not generalize to fields with different contribution recording methods.
@@ VII Conclusion
VIII Acknowledgement
This research is supported in part by the Natural Sciences and Engineering Research Council of Canada (NSERC) Discovery Grants program, the Canada Foundation for Innovation’s John R. Evans Leaders Fund (CFI-JELF), and by the industry-stream NSERC CREATE in Software Analytics Research (SOAR).